A fire department in a rural county reports that its response time to fires is approximately
Normally distributed with a mean of 22 minutes and a standard deviation of 11.9 minutes.
Approximately what proportion of their response times is over 30 minutes?
a)0.03
b)0.21
c)0.25
d)0.75
e)0.79

Answer by ewatrrr(24785)

 
Hi,Re TY
You did not say how You determine the P-value given the z-value
NORMSDIST(.6723) is one such way of doing that ...using an Excel function NORMSDIST.
Chart or TI calculator works as well.
μ = 22 , σ =11.9 and z = (30-22)/11.9
P(x>30) = 1 -P(z ≤ .6723) = 1- NORMSDIST(.6723)= 1-.75 = .25